# Harder-Narasimhan filtrations of decorated vector bundles

### v1

@@ 2. Decorated vector bundles
Recall that a decorated vector bundle $(E,s)$ is a vector bundle $E$ equipped with a reduction $s$ to a complex reductive subgroup $G\subseteq\mathbf{GL}(r;\mathbb{C})$. Examples include symplectic and special-orthogonal vector bundles, as well as vector bundles with trivial determinants, i.e., special vector bundles. As we did for special vector bundles in Theorem 1.12, we recall conditions for symplectic and special-orthogonal vector bundles to be Ramanathan-(semi)-stable in Theorems 2.6 and 2.13.
@@ 2.1. Symplectic vector bundles
Both variants are isomorphic by conjugation in $\mathbf{GL}(2n,\mathbb{C})$. Our variant allows us to intersect our earlier choices of Cartan and Borel subgroups of $\mathbf{GL}(2n,\mathbb{C})$ to obtain Cartan and Borel subgroups of $\mathbf{Sp}(2n,\mathbb{C})$. This has the advantage that all standard parabolic subgroups of $\mathbf{Sp}(2n,\mathbb{C})$ appear as upper-right block matrices, allowing us to more easily give symplectic versions of Lemmas 1.3 and 1.10. However, the root system of $\mathbf{Sp}(2n,\mathbb{C})$ is quite different to that of $\mathbf{GL}(2n,\mathbb{C})$ and $\mathbf{SL}(2n,\mathbb{C})$. Standard parabolic subgroups of $\mathbf{Sp}(2n,\mathbb{C})$ are in correspondence with subsets $I$ of the set of simple roots $\triangle$:
@@
For a symplectic vector bundle $(E,\beta)$, a symplectic frame bundle $\mathrm{Fr}_{\mathbf{Sp}}(E,\beta)$ is a principal-$\mathbf{Sp}(2n,\mathbb{C})$-bundle, whose fiber at $x\in X$ is given by:
@@
We obtain a symplectic variant of Lemmas 1.3 and 1.6.
@@
For the case of $F=F^{\perp}$, the calculations are the same, apart from removing the middle rows and columns in (2.5,2.6), so the isomorphism in (2.4) follows. ∎
@@
For (i) to (ii), the inclusion $\mathbf{Sp}(2n,\mathbb{C})\hookrightarrow\mathbf{GL}(2n,\mathbb{C})$ maps $\mathbf{R}(\mathbf{Sp}(2n,\mathbb{C}))=0$ into $\mathbf{R}(\mathbf{GL}(2n,\mathbb{C}))$. This extends $\mathrm{Fr}_{\mathbf{Sp}}(E,\beta)$ as a principal-$\mathbf{Sp}(2n,\mathbb{C})$-bundle to $\mathrm{Fr}(E)$ as a principal-$\mathbf{GL}(2n,\mathbb{C})$-bundle, so $\mathrm{Fr}(E)$ is Ramanathan-semistable due to Theorem 1.14. Then $E$ is slope-semistable following Theorem 1.11.
@@ 2.2. Special-orthogonal vector bundles
Much like for symplectic vector bundles, Ramanathan-(semi)-stability corresponds to checking slope conditions on isotropic subbundles, defined analogously to Definition 2.1. To see this, we need to understand the root system of the special-orthogonal group $\mathbf{SO}(r,\mathbb{C})$, which we define with respect to the nondegenerate symmetric form $\beta(x,y):=\sum_{i=1}^{r}x_{r+1-i}y_{i}$:
@@
using notation from (1.1,2.2). This difference in parity is significant, as it affects how standard parabolic subgroups appear. In the odd case, standard parabolic subgroups are all subgroups of upper-right block matrices. In the even case, not all standard parabolic subgroups are upper-right block matrices, as seen concretely in the following example.
@@
Using the fact that parabolic subgroups always arise as stabilizers of flags, as seen in [conrad_standardparabolicsubgroups, Theorem 3.9 (i)], we can use [conrad_standardparabolicsubgroups, Corollary 3.4] to at least say that in the even case, parabolic subgroups of $\mathbf{SO}(r,\mathbb{C})$ are conjugate, via an element $g\in\mathbf{O}(r,\mathbb{C})$, to upper-right block matrix subgroups of $\mathbf{SO}(r,\mathbb{C})$. Now we can proceed as in the symplectic case, being careful to respect odd and even cases. We define special-orthogonal frame bundles, similarly to Definitions 1.5 and 2.3.
@@
For a special-orthogonal vector bundle $(E,\beta,\tau)$, the special-orthogonal frame bundle $\mathrm{Fr}_{\mathbf{SO}}(E,\beta,\tau)$ is a principal-$\mathbf{SO}(r,\mathbb{C})$-bundle, whose fiber at $x\in X$ is given by:
@@
We obtain a special-orthogonal variant of Lemmas 1.3, 1.6, and 2.4. Let $(E,\beta,\tau)$ be a special-orthogonal vector bundle of rank $r\geqslant 3$.
@@
Lemma 2.11 leads to a special-orthogonal version of Lemmas 1.10 and 2.5.
@@
where $(\alpha_{i,j})_{i,j\in J}$ are cocycles of $F$ and $\beta_{i,j}(x)\in\mathbb{C}^{\times}$. Since the cocycles $(\sigma_{i,j})_{i,j\in J}$ map into $P_{\{\alpha_{n}+\alpha_{n+1}\}}$, they define a reduction of $\mathrm{Fr}_{\mathbf{SO}}(E,\beta,\tau)$ to $P_{\{\alpha_{n}+\alpha_{n+1}\}}$, corresponding to a Lagrangian subbundle $F^{\prime}$ of $E$ strictly containing $F$. At every fiber $E_{x}$, $x\in X$, the Lagrangian subspace $F^{\prime}_{x}$ strictly containing $F_{x}$ exists following [conrad_standardparabolicsubgroups, Theorem 3.9 (ii)]. This is a quirk of the specific reductive group structure of $\mathbf{SO}(2n,\mathbb{C})$, where parabolic subgroups are not in bijection with isotropic flags of $\mathbb{C}^{2n}$. Much like in Example 2.9, we can conjugate the cocycles $(\sigma_{i,j})_{i,j\in J}$ by an orthogonal matrix swapping the middle rows and columns, allowing us to swap $(\beta_{i,j})_{i,j\in J}$ with $(\beta_{i,j}^{-1})_{i,j\in J}$. In this way, we can assume $\deg(F^{\prime}/F)\geqslant 0$ without loss of generality. Since $\deg(F)+\deg(F^{\prime}/F)=\deg(F^{\prime})(\leqslant)0$, and since $\deg(F^{\prime}/F)\geqslant 0$, we have $\deg(F)(\leqslant)0$, and so $\mu(F)(\leqslant)\mu(E)=0$. ∎
@@ 3.1. Harder-Narasimhan filtrations of decorated vector bundles
We wish to construct Harder-Narasimhan filtrations of symplectic and special-orthogonal vector bundles. Such filtrations of vector bundles were first constructed by Harder and Narasimhan in [harder-narasimhan_onthecohomologygroupsofmodulispaces, Lemma 1.3.8].
@@
This filtration measures how far away the vector bundle $E$ is from being slope-semistable, as $E$ is slope-semistable if and only if its Harder-Narasimhan filtration is trivial. As we will see in the proof, each quotient $F_{m}$ contradicts the slope-semistability of a quotient bundle of $E$, so the Harder-Narasimhan filtration measures to what extent $E$ is slope-semistable. There exist analogous filtrations for symplectic and special-orthogonal vector bundles. As seen in [atiyah-bott_yangmills, $\S$10.] and [biswas-holla_harder-narasimhanreductionofprincbund], these filtrations come from canonical reductions of principal bundles. We sketch a proof of Theorem 3.1 by constructing the filtration by hand.
@@
For our purposes, it is important to know just the main steps of the proof, so we refer to [harder-narasimhan_onthecohomologygroupsofmodulispaces, Lemma 1.3.8] for further details.
@@
which we claim are bounded. We first handle $\deg_{\max}(E)$. The 0-th sheaf cohomology of a vector bundle is the complex vector space of global sections. For a subbundle $F$ of $E$ of rank $l=1,\ldots,r$, a global section of $F$ is also a global section of $E$, and thus $\dim_{\mathbb{C}}(H^{0}(X,E))\geqslant\dim_{\mathbb{C}}(H^{0}(X,F))$. We denote the genus of $X$ by $g$, we then have due to Riemann-Roch, from [*][2. Theorem 2]friedman_algebraicsurfacesandholomorphicvectorbundles, that:
@@
Verify that $F$ is the strongly contradicting semistability (SCSS) subbundle $\mathcal{G}(E)$ of $E$, as defined in [harder-narasimhan_onthecohomologygroupsofmodulispaces, 1.3], and is hence unique, as shown in [harder-narasimhan_onthecohomologygroupsofmodulispaces, Proposition 1.3.4]. Being SCCS means $\mathcal{G}(E)$ is slope-semistable, and if $\mathcal{G}(E)\subsetneq E$, then $\mu(\mathcal{G}(E))>\mu_{\max}(E/\mathcal{G}(E))$. One way to prove uniqueness is by showing that the projection $\mathcal{G}(E)\rightarrow E/\mathcal{G}(E)^{\prime}$ to another SCSS bundle $\mathcal{G}(E)^{\prime}$ is trivial, i.e., equal to 0, using that $\mathcal{G}(E)$ is slope-semistable, and $\mu(\mathcal{G}(E))>\mu_{\max}(E/\mathcal{G}(E)^{\prime})$.
@@
is the unique Harder-Narasimhan filtration of $E/E_{1}$. It remains to show $E_{1}^{\prime}=E_{1}$, which follows from the uniqueness of SCSS bundles in [harder-narasimhan_onthecohomologygroupsofmodulispaces, Proposition 1.3.4]. ∎
Special vector bundles and their underlying vector bundles have the same slope-semistability conditions, coming from the Ramanathan-semistability of their special (frame) bundles, following Theorems 1.11 and 1.12, and Lemma 1.13. Thus, constructing Harder-Narasimhan filtrations for special vector bundles is identical to Theorem 3.1.
@@
is the unique symplectic Harder-Narasimhan filtration of $(E_{1}^{\perp}/E_{1},\beta)$. It remains to show $E_{1}^{\prime}=E_{1}$, which is done in multiple steps: First, the projection $E_{1}\rightarrow E/E_{1}^{\prime\perp}$ is trivial, i.e., equal to 0, using that $E_{1}$ is slope-semistable, and $\mu(E_{1})>0>\mu_{\max}(E/E_{1}^{\prime\perp})=\mu(E/E_{1}^{\prime\perp})$. The second inequality is true since $E/E_{1}^{\prime\perp}$ is isomorphic to the dual bundle $E_{1}^{\prime*}$ of $E_{1}^{\prime}$, where $E_{1}^{\prime*}$ is slope-semistable, and where $\mu(E_{1}^{\prime})>0$ implies $0>\mu(E_{1}^{\prime*})$. Next, the projection $E_{1}\rightarrow E_{1}^{\prime\perp}/E_{2}^{\prime\perp}$ is trivial for similar reasons, since $E_{1}$ is slope semistable, and $\mu(E_{1})>0>\mu_{\max}(E_{1}^{\prime\perp}/E_{2}^{\prime\perp})=\mu(E_{1}^{\prime\perp}/E_{2}^{\prime\perp})$. Recursively, the projection $E_{1}\rightarrow E_{m-1}^{\prime\perp}/E_{m}^{\prime\perp}$ is trivial for all $m=1,\ldots,k$, and thus $E_{1}\subseteq E_{t}^{\prime\perp}$. If $E_{t}^{\prime}$ is Lagrangian, then the projection $E_{1}\rightarrow E_{t}^{\prime\perp}/E_{t}^{\prime}$ is trivial, since $E_{t}^{\prime\perp}/E_{t}^{\prime}=0$. Otherwise, $E_{t}^{\prime}$ is not Lagrangian, and to prove that $E_{1}\rightarrow E_{t}^{\prime\perp}/E_{t}^{\prime}$ is trivial, we use that $E_{1}$ is slope-semistable, and $\mu(E_{1})>0=\mu_{\max}(E_{t}^{\prime\perp}/E_{t}^{\prime})=\mu(E_{t}^{\prime\perp}/E_{t}^{\prime})$. The last equality follows from Theorem 2.6 and Lemma 2.7, and the fact that $\mu(E_{t}^{\prime\perp}/E_{t}^{\prime})=0$. Due to $E_{1}\subseteq E_{t}^{\prime}$, both $E_{1}$ and $E_{1}^{\prime}$ are SCCS subbundles of $E_{t}^{\prime}$, which are unique due to [harder-narasimhan_onthecohomologygroupsofmodulispaces, Proposition 1.3.4]. Hence, $E_{1}=E_{1}^{\prime}$. ∎
@@
fulfilling (i), (ii), and (iii), which contradicts the uniqueness of Harder-Narasimhan filtrations. This extension was mentioned in the proof of Theorem 2.13, where we constructed an isotropic subbundle $F^{\prime}$ containing an isotropic subbundle $F=E_{t}$ of rank n1n-1, using [conrad_standardparabolicsubgroups, Theorem 3.9 (ii)] at every fiber $F_{x}$ of $F$. From the reductive groups we study, this is only possible in $\mathbf{SO}(2n,\mathbb{C})$, where parabolic subgroups are not in bijection with isotropic flags of $\mathbb{C}^{2n}$.
@@
By comparing all the variants of Harder-Narasimhan filtrations from Theorems 3.1-3.5, it is clear that the symplectic and special-orthogonal Harder-Narasimhan filtrations:
@@
For a principal-$G$-bundle $\xi$, where $G$ is complex and reductive, there exists a special-orthogonal vector bundle structure on the adjoint bundle $\mathrm{ad}(\xi)$.
@@
Since $G$ is reductive, there exists a nondegenerate adjoint-invariant form on $\mathrm{Lie}(G)$. Using that $\mathrm{ad}(\xi)$ is induced by the extension of $\xi$ by the adjoint representation $\mathrm{Ad}:G\rightarrow\mathbf{GL}(\mathrm{Lie}(G))$, this form induces a nondegenerate symmetric form $\beta$ on $\mathrm{ad}(\xi)$, such that $(\mathrm{ad}(\xi),\beta)$ is an orthogonal vector bundle. To show that $(\mathrm{ad}(\xi),\beta)$ is special-orthogonal, we must show that the determinant bundle $\det(\mathrm{ad}(\xi))$ is isomorphic to the product bundle $X\times\mathbb{C}$. Since $G$ is an affine group, it can be viewed as a matrix subgroup of $\mathbf{GL}(r,\mathbb{C})$, following [borel_linearalgebraicgroups, I. 1.10 Proposition], so the adjoint representation $\mathrm{Ad}:G\rightarrow\mathbf{GL}(\mathrm{Lie}(G))$ can be viewed as matrix conjugation within $\mathbf{GL}(r,\mathbb{C})$. Cocycles $(\sigma_{i,j})_{i,j\in J}$ of $\xi$ induce cocycles $(\mathrm{Ad}\circ\sigma_{i,j})_{i,j\in J}$ of $\mathrm{ad}(\xi)$, equal to $((\sigma_{i,j}^{-1})^{T}\otimes\sigma_{i,j})_{i,j\in J}$. This induces cocycles $(\det((\sigma_{i,j}^{-1})^{T}\otimes\sigma_{i,j}))_{i,j\in J}$ of $\det(\mathrm{ad}(\xi))$, which are cocycles of the product bundle $X\times\mathbb{C}$. Thus, the claim follows. ∎
Lemma 3.7 does not give a unique special-orthogonal vector bundle structure on $\mathrm{ad}(\xi)$, as we made the choices of an adjoint-invariant form on $\mathrm{Lie}(G)$, cocycles $(\sigma_{i,j})_{i,j\in J}$ of $\xi$, and an embedding of $G$ into $\mathbf{GL}(r,\mathbb{C})$. However, in concrete examples, e.g. $G=\mathbf{Sp}(2n,\mathbb{C})$ or $G=\mathbf{SO}(r,\mathbb{C})$, these are often canonical. For the upcoming example of special-orthogonal Harder-Narasimhan filtrations (Example 3.9), we will make use of the following result from [*][Lemma 10.1]atiyah-bott_yangmills.
@@
For the slope-semistability of $E\otimes E^{\prime}$, we first assume that $E$ and $E^{\prime}$ are slope-stable. In [wells-garciaprada_differentialanalysis, Appendix 4.1-4.2], we see that subbundles $E$ and $E^{\prime}$ correspond to irreducible unitary representations of an extension $\Gamma$ of $\pi_{1}(X)$. By tensoring these representations, we get another unitary representation of $\Gamma$ that corresponds to $E\otimes E^{\prime}$, which is thus slope-semistable. Following [*][Lemma 10.1]atiyah-bott_yangmills, to generalize to the situation where $E$ and $E^{\prime}$ are only slope-semistable, we may use Jordan-Hölder filtrations of $E$ and $E^{\prime}$, as discussed in [huybrechts-lehn_geometryofmodulispaces, 1.5]. ∎
With Lemmas 3.7 and 3.8, we can construct an example of a special-orthogonal vector bundle, and find its special-orthogonal Harder-Narasimhan filtration.

### v2

@@ 3. Decorated vector bundles
Recall that a decorated vector bundle $(E,s)$ is a vector bundle $E$ equipped with a reduction $s$ to a complex reductive subgroup $G\subseteq\mathbf{GL}(r;\mathbb{C})$. Examples include symplectic and special-orthogonal vector bundles, as well as vector bundles with trivial determinants, i.e., special vector bundles. As we did for special vector bundles in Theorem 2.12, we recall conditions for symplectic and special-orthogonal vector bundles to be Ramanathan-(semi)-stable in Theorems 3.6 and 3.13.
@@ 3.1. Symplectic vector bundles
Both variants are isomorphic by conjugation in $\mathbf{GL}(2n,\mathbb{C})$. Our variant allows us to intersect our earlier choices of Cartan and Borel subgroups of $\mathbf{GL}(2n,\mathbb{C})$ to obtain Cartan and Borel subgroups of $\mathbf{Sp}(2n,\mathbb{C})$. This has the advantage that all standard parabolic subgroups of $\mathbf{Sp}(2n,\mathbb{C})$ appear as upper-right block matrices, allowing us to more easily give symplectic versions of Lemmas 2.3 and 2.10. However, the root system of $\mathbf{Sp}(2n,\mathbb{C})$ is quite different to that of $\mathbf{GL}(2n,\mathbb{C})$ and $\mathbf{SL}(2n,\mathbb{C})$. Standard parabolic subgroups of $\mathbf{Sp}(2n,\mathbb{C})$ are in correspondence with subsets $I$ of the set of simple roots $\triangle$:
@@
For a symplectic vector bundle $(E,\beta)$, a symplectic frame bundle $\mathrm{Fr}_{\mathbf{Sp}}(E,\beta)$ is a principal $\mathbf{Sp}(2n,\mathbb{C})$-bundle, whose fiber at $x\in X$ is given by:
@@
We obtain a symplectic variant of Lemmas 2.3 and 2.6.
@@
For the case of $F=F^{\perp}$, the calculations are the same, apart from removing the middle rows and columns in (3.5,3.6), so the isomorphism in (3.4) follows. ∎
@@
For (i) to (ii), the inclusion $\mathbf{Sp}(2n,\mathbb{C})\hookrightarrow\mathbf{GL}(2n,\mathbb{C})$ maps $\mathbf{R}(\mathbf{Sp}(2n,\mathbb{C}))=0$ into $\mathbf{R}(\mathbf{GL}(2n,\mathbb{C}))$. This extends $\mathrm{Fr}_{\mathbf{Sp}}(E,\beta)$ as a principal $\mathbf{Sp}(2n,\mathbb{C})$-bundle to $\mathrm{Fr}(E)$ as a principal $\mathbf{GL}(2n,\mathbb{C})$-bundle, so $\mathrm{Fr}(E)$ is Ramanathan-semistable due to Theorem 2.14. Then $E$ is slope-semistable following Theorem 2.11.
@@ 3.2. Special-orthogonal vector bundles
Much like for symplectic vector bundles, Ramanathan-(semi)-stability corresponds to checking slope conditions on isotropic subbundles, as seen in [Ram81, $\S$4]. To see this, we need to understand the root system of the special-orthogonal group $\mathbf{SO}(r,\mathbb{C})$, which we define with respect to the nondegenerate symmetric form $\beta(x,y):=\sum_{i=1}^{r}x_{r+1-i}y_{i}$:
@@
using notation from (2.1,3.2). This difference in parity is significant, as it affects how standard parabolic subgroups appear. In the odd case, standard parabolic subgroups are all subgroups of upper-right block matrices. In the even case, not all standard parabolic subgroups are upper-right block matrices, as seen concretely in the following example.
@@
Using the fact that parabolic subgroups always arise as stabilizers of flags, as seen in [Con16, Theorem 3.9 (i)], we can use [Con16, Corollary 3.4] to at least say that in the even case, parabolic subgroups of $\mathbf{SO}(r,\mathbb{C})$ are conjugate, via an element $g\in\mathbf{O}(r,\mathbb{C})$, to upper-right block matrix subgroups of $\mathbf{SO}(r,\mathbb{C})$. Now we can proceed as in the symplectic case, being careful to respect odd and even cases. We define special-orthogonal frame bundles, similarly to Definitions 2.5 and 3.3.
@@
For a special-orthogonal vector bundle $(E,\beta,\tau)$, the special-orthogonal frame bundle $\mathrm{Fr}_{\mathbf{SO}}(E,\beta,\tau)$ is a principal $\mathbf{SO}(r,\mathbb{C})$-bundle, whose fiber at $x\in X$ is given by:
@@
We obtain a special-orthogonal variant of Lemmas 2.3, 2.6, and 3.4. Let $(E,\beta,\tau)$ be a special-orthogonal vector bundle of rank $r\geqslant 3$.
@@
Lemma 3.11 leads to a special-orthogonal version of Lemmas 2.10 and 3.5.
@@
where $(\alpha_{i,j})_{i,j\in J}$ are cocycles of $F$ and $\beta_{i,j}(x)\in\mathbb{C}^{\times}$. Since the cocycles $(\sigma_{i,j})_{i,j\in J}$ map into $P_{\{\alpha_{n}+\alpha_{n+1}\}}$, they define a reduction of $\mathrm{Fr}_{\mathbf{SO}}(E,\beta,\tau)$ to $P_{\{\alpha_{n}+\alpha_{n+1}\}}$, corresponding to a Lagrangian subbundle $F^{\prime}$ of $E$ strictly containing $F$. At every fiber $E_{x}$, $x\in X$, the Lagrangian subspace $F^{\prime}_{x}$ strictly containing $F_{x}$ exists following [Con16, Theorem 3.9 (ii)]. This is a quirk of the specific reductive group structure of $\mathbf{SO}(2n,\mathbb{C})$, where parabolic subgroups are not in bijection with isotropic flags of $\mathbb{C}^{2n}$. Much like in Example 3.9, we can conjugate the cocycles $(\sigma_{i,j})_{i,j\in J}$ by an orthogonal matrix swapping the middle rows and columns, allowing us to swap $(\beta_{i,j})_{i,j\in J}$ with $(\beta_{i,j}^{-1})_{i,j\in J}$. In this way, we can assume $\deg(F^{\prime}/F)\geqslant 0$ without loss of generality. Since $\deg(F)+\deg(F^{\prime}/F)=\deg(F^{\prime})(\leqslant)0$, and since $\deg(F^{\prime}/F)\geqslant 0$, we have $\deg(F)(\leqslant)0$, and so $\mu(F)(\leqslant)\mu(E)=0$. ∎
@@ 4.1. Harder-Narasimhan filtrations of decorated vector bundles
We wish to construct Harder-Narasimhan filtrations of symplectic and special-orthogonal vector bundles. Such filtrations of vector bundles were first constructed by Harder and Narasimhan in [HN75, Lemma 1.3.8].
@@
This filtration measures how far away the vector bundle $E$ is from being slope-semistable, as $E$ is slope-semistable if and only if its Harder-Narasimhan filtration is trivial. As we will see in the proof, each quotient $F_{m}$ contradicts the slope-semistability of a quotient bundle of $E$, so the Harder-Narasimhan filtration measures to what extent $E$ is slope-semistable. There exist analogous filtrations for symplectic and special-orthogonal vector bundles. As seen in [AB83, $\S$10] and [BH04], these filtrations come from canonical reductions of principal bundles. We sketch a proof of Theorem 4.1 by constructing the filtration by hand.
@@
For our purposes, it is important to know just the main steps of the proof, so we refer to [HN75, Lemma 1.3.8] for further details.
@@
which we claim are bounded. We first handle $\deg_{\max}(E)$. The 0-th sheaf cohomology of a vector bundle is the complex vector space of global sections. For a subbundle $F$ of $E$ of rank $l=1,\ldots,r$, a global section of $F$ is also a global section of $E$, and thus $\dim_{\mathbb{C}}(H^{0}(X,E))\geqslant\dim_{\mathbb{C}}(H^{0}(X,F))$. We denote the genus of $X$ by $g$, we then have due to Riemann-Roch, from [Fri98, 2. Theorem 2], that:
@@
Verify that $F$ is the strongly contradicting semistability (SCSS) subbundle $\mathcal{G}(E)$ of $E$, as defined in [HN75, 1.3], and is hence unique, as shown in [HN75, Proposition 1.3.4]. Being SCCS means $\mathcal{G}(E)$ is slope-semistable, and if $\mathcal{G}(E)\subsetneq E$, then $\mu(\mathcal{G}(E))>\mu_{\max}(E/\mathcal{G}(E))$. One way to prove uniqueness is by showing that the projection $\mathcal{G}(E)\rightarrow E/\mathcal{G}(E)^{\prime}$ to another SCSS bundle $\mathcal{G}(E)^{\prime}$ is trivial, i.e., equal to 0, using that $\mathcal{G}(E)$ is slope-semistable, and $\mu(\mathcal{G}(E))>\mu_{\max}(E/\mathcal{G}(E)^{\prime})$.
@@
is the unique Harder-Narasimhan filtration of $E/E_{1}$. It remains to show $E_{1}^{\prime}=E_{1}$, which follows from the uniqueness of SCSS bundles in [HN75, Proposition 1.3.4]. ∎
Special vector bundles and their underlying vector bundles have the same slope-semistability conditions, coming from the Ramanathan-semistability of their special (frame) bundles, following Theorems 2.11 and 2.12, and Lemma 2.13. Thus, constructing Harder-Narasimhan filtrations for special vector bundles is identical to Theorem 4.1.
@@
is the unique symplectic Harder-Narasimhan filtration of $(E_{1}^{\perp}/E_{1},\beta)$. It remains to show $E_{1}^{\prime}=E_{1}$, which is done in multiple steps: First, the projection $E_{1}\rightarrow E/E_{1}^{\prime\perp}$ is trivial, i.e., equal to 0, using that $E_{1}$ is slope-semistable, and $\mu(E_{1})>0>\mu_{\max}(E/E_{1}^{\prime\perp})=\mu(E/E_{1}^{\prime\perp})$. The second inequality is true since $E/E_{1}^{\prime\perp}$ is isomorphic to the dual bundle $E_{1}^{\prime*}$ of $E_{1}^{\prime}$, where $E_{1}^{\prime*}$ is slope-semistable, and where $\mu(E_{1}^{\prime})>0$ implies $0>\mu(E_{1}^{\prime*})$. Next, the projection $E_{1}\rightarrow E_{1}^{\prime\perp}/E_{2}^{\prime\perp}$ is trivial for similar reasons, since $E_{1}$ is slope-semistable, and $\mu(E_{1})>0>\mu_{\max}(E_{1}^{\prime\perp}/E_{2}^{\prime\perp})=\mu(E_{1}^{\prime\perp}/E_{2}^{\prime\perp})$. Recursively, the projection $E_{1}\rightarrow E_{m-1}^{\prime\perp}/E_{m}^{\prime\perp}$ is trivial for all $m=1,\ldots,k$, and thus $E_{1}\subseteq E_{t}^{\prime\perp}$. If $E_{t}^{\prime}$ is Lagrangian, then the projection $E_{1}\rightarrow E_{t}^{\prime\perp}/E_{t}^{\prime}$ is trivial, since $E_{t}^{\prime\perp}/E_{t}^{\prime}=0$. Otherwise, $E_{t}^{\prime}$ is not Lagrangian, and to prove that $E_{1}\rightarrow E_{t}^{\prime\perp}/E_{t}^{\prime}$ is trivial, we use that $E_{1}$ is slope-semistable, and $\mu(E_{1})>0=\mu_{\max}(E_{t}^{\prime\perp}/E_{t}^{\prime})=\mu(E_{t}^{\prime\perp}/E_{t}^{\prime})$. The last equality follows from Theorem 3.6 and Lemma 3.7, and the fact that $\mu(E_{t}^{\prime\perp}/E_{t}^{\prime})=0$. Due to $E_{1}\subseteq E_{t}^{\prime}$, both $E_{1}$ and $E_{1}^{\prime}$ are SCCS subbundles of $E_{t}^{\prime}$, which are unique due to [HN75, Proposition 1.3.4]. Hence, $E_{1}=E_{1}^{\prime}$. ∎
@@
fulfilling (i), (ii), and (iii), which contradicts the uniqueness of Harder-Narasimhan filtrations. This extension was mentioned in the proof of Theorem 3.13, where we constructed an isotropic subbundle $F^{\prime}$ containing an isotropic subbundle $F=E_{t}$ of rank n1n-1, using [Con16, Theorem 3.9 (ii)] at every fiber $F_{x}$ of $F$. From the reductive groups we study, this is only possible in $\mathbf{SO}(2n,\mathbb{C})$, where parabolic subgroups are not in bijection with isotropic flags of $\mathbb{C}^{2n}$.
@@
By comparing all the variants of Harder-Narasimhan filtrations from Theorems 4.1-4.5, it is clear that the symplectic and special-orthogonal Harder-Narasimhan filtrations:
@@
For a principal $G$-bundle $\xi$, where $G$ is complex and reductive, there exists a special-orthogonal vector bundle structure on the adjoint bundle $\mathrm{ad}(\xi)$.
@@
Since $G$ is reductive, there exists a nondegenerate adjoint-invariant form on $\mathrm{Lie}(G)$. Using that $\mathrm{ad}(\xi)$ is induced by the extension of $\xi$ by the adjoint representation $\mathrm{Ad}:G\rightarrow\mathbf{GL}(\mathrm{Lie}(G))$, this form induces a nondegenerate symmetric form $\beta$ on $\mathrm{ad}(\xi)$, such that $(\mathrm{ad}(\xi),\beta)$ is an orthogonal vector bundle. To show that $(\mathrm{ad}(\xi),\beta)$ is special-orthogonal, we must show that the determinant bundle $\det(\mathrm{ad}(\xi))$ is isomorphic to the product bundle $X\times\mathbb{C}$. Since $G$ is an affine group, it can be viewed as a matrix subgroup of $\mathbf{GL}(r,\mathbb{C})$, following [Bor91, I. 1.10 Proposition], so the adjoint representation $\mathrm{Ad}:G\rightarrow\mathbf{GL}(\mathrm{Lie}(G))$ can be viewed as matrix conjugation within $\mathbf{GL}(r,\mathbb{C})$. Cocycles $(\sigma_{i,j})_{i,j\in J}$ of $\xi$ induce cocycles $(\mathrm{Ad}\circ\sigma_{i,j})_{i,j\in J}$ of $\mathrm{ad}(\xi)$, equal to $((\sigma_{i,j}^{-1})^{T}\otimes\sigma_{i,j})_{i,j\in J}$. This induces cocycles $(\det((\sigma_{i,j}^{-1})^{T}\otimes\sigma_{i,j}))_{i,j\in J}$ of $\det(\mathrm{ad}(\xi))$, which are cocycles of the product bundle $X\times\mathbb{C}$. Thus, the claim follows. ∎
Lemma 4.7 does not give a unique special-orthogonal vector bundle structure on $\mathrm{ad}(\xi)$, as we made the choices of an adjoint-invariant form on $\mathrm{Lie}(G)$, cocycles $(\sigma_{i,j})_{i,j\in J}$ of $\xi$, and an embedding of $G$ into $\mathbf{GL}(r,\mathbb{C})$. However, in concrete examples, e.g. $G=\mathbf{Sp}(2n,\mathbb{C})$ or $G=\mathbf{SO}(r,\mathbb{C})$, these are often canonical. For the upcoming example of special-orthogonal Harder-Narasimhan filtrations (Example 4.9), we will make use of the following result from [AB83, Lemma 10.1].
@@
For the slope-semistability of $E\otimes E^{\prime}$, we first assume that $E$ and $E^{\prime}$ are slope-stable. In [WGP08, Appendix 4.1-4.2], we see that subbundles $E$ and $E^{\prime}$ correspond to irreducible unitary representations of an extension $\Gamma$ of $\pi_{1}(X)$. By tensoring these representations, we get another unitary representation of $\Gamma$ that corresponds to $E\otimes E^{\prime}$, which is thus slope-semistable. Following [AB83, Lemma 10.1], to generalize to the situation where $E$ and $E^{\prime}$ are only slope-semistable, we may use Jordan-Hölder filtrations of $E$ and $E^{\prime}$, as discussed in [HL97, 1.5]. ∎
With Lemmas 4.7 and 4.8, we can construct an example of a special-orthogonal vector bundle, and find its special-orthogonal Harder-Narasimhan filtration.